# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v1

@@ 4.2 Few-Shot Fine-Tuning Experiments
Although AlphaDE does not utilize fitness labels of protein sequences in the pretrain, fine-tune, or MCTS inference, we additionally test whether the protein fitness influences the performance of AlphaDE by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3 and are very similar to Figure 2, showing that AlphaDE’s evolution capability is unaffected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly fitness measurement assays. We also show that AlphaDE works with the sequences from homology searching (potter_hmmer_2018) in Appendix J.
@@ 4.4 Condensing Protein Sequence Space
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (abramson_accurate_2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
Task AlphaDE TreeNeuralTS TreeNeuralUCB EvoPlay PEX
avGFP 3.991 3.707 3.785 3.835 3.754
UBE2I 3.989 3.632 3.653 0.647 3.101

### v2

@@ 4.2 Few-Shot Fine-Tuning Experiments
Although AlphaDE does not utilize fitness labels of protein sequences in the pretrain, fine-tune, or MCTS inference, we additionally test whether the protein fitness influences the performance of AlphaDE by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3 and are very similar to Figure 2, showing that AlphaDE’s evolution capability is unaffected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly fitness measurement assays. We also show that AlphaDE works with the sequences from homology searching such as HMMER (potter_hmmer_2018) in Appendix J.
@@ 4.4 Condensing Protein Sequence Space
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (abramson_accurate_2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the comparison of AlphaDE without the fine-tuning step and the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
Task AlphaDE MLDE TreeNeuralTS TreeNeuralUCB EvoPlay PEX
avGFP 3.991 2.443 3.707 3.785 3.835 3.754
UBE2I 3.989 1.434 3.632 3.653 0.647 3.101
@@ Appendix I Diversity of Evolved Sequences by AlphaDE
At the same time, we also compare with AlphaDE with pretrained protein language model to validate whether it recovers the folded structure of wild-type avGFP. We follow the same pipeline as condensing avGFP with standard AlphaDE. The comparison of the AlphaDE with pretrained ESM2-35M and standard AlphaDE is shown in Table 14 below. It is clear that the condensed avGFPs by AlphaDE with pretrained ESM2-35M have much larger RMSDs, indicating the recovered folded structures are less similar to the wild-type avGFP structure than the standard AlphaDE.
Table 14: Comparison of AlphaDE with pretrained ESM2-35M and standard AlphaDE with fine-tuned ESM2-35M for condensing avGFP. The metric is RMSD to the wild-type avGFP (PDB ID: 1EMA).
RMSD ($\downarrow$) 1 2 3 4 5
standard AlphaDE with fine-tuned ESM2-35M 1.10 1.40 1.70 1.83 1.92
AlphaDE with pretrained ESM2-35M 5.92 6.27 6.79 7.55 8.08